OpenStudy (anonymous):

I REALLY NEED HELP WITH ALGEBRA!!! I don't understand these problems that well: 1. -2(x-3) = 2(x+1) 2. 2x + 4 = 4(x - 5) Can someone please explain how to do at least one of them step by step?! Thanks

OpenStudy (anonymous):

are you solving for x in these?

OpenStudy (anonymous):

helloo?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok so for the 1st one multiply -2 and 2 in the equation. so (-2 times x) (-2 time -3) = (2 times x) (2 times 1)

OpenStudy (anonymous):

like multiply -2 and 3 on the left side and multiply 2 and 1 on the right side?

OpenStudy (anonymous):

yeah and tell me what you get

OpenStudy (anonymous):

also is this problem in a book? if yes what is it called

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I get -6 and 2

OpenStudy (anonymous):

i got* lol

OpenStudy (anonymous):

after doing what i said you should of gotten -2x+6=2x+1 idk if im right eh. could you give me the title or name of the chapter or page?

OpenStudy (anonymous):

after getting -2x+6=2x+1 you solve for x from there

OpenStudy (anonymous):

not 100% sure if correct

OpenStudy (anonymous):

well yeah i just thought you meant specifically those numbers but i know how to do the first step (i did mess up on the -2 times -3 tho) What's the next step?

OpenStudy (anonymous):

6-1

OpenStudy (anonymous):

so it would be -2x +5=2x is that right?

OpenStudy (anonymous):

yeah you know what to do next?

OpenStudy (anonymous):

um.. would i do -2x +5=2x -5 -5 or -2x +5=2x +2x +2x

OpenStudy (anonymous):

you would do the add 2x one because you cant add subtract unlike terms. those two go together because they both have x

OpenStudy (anonymous):

That's kinda what i thought lol. -2x +5=2x +2x +2x 5=4x would x= 1.25? or did i do something wrong?

OpenStudy (anonymous):

i got that too. lol thats the thing in unsure on. when i typed in the equation here https://mathway.com/ they said that it was 1

OpenStudy (anonymous):

> n < it would kinda help if i knew what that equation was called or something

OpenStudy (anonymous):

oh. Thats weird.. I'm gonna try to do it again and see if i get 1 or 1.25

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

is this algebra 1 or 2 stuff?

OpenStudy (anonymous):

if you see me typing it just me trying to solve it & Its algebra 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you do have a algebra 1 book right?

OpenStudy (anonymous):

Not exactly.. I'm doing online school this year

OpenStudy (anonymous):

ohh ok then you know what the thing ur working on is called right?

OpenStudy (anonymous):

One step equations

OpenStudy (anonymous):

no

OpenStudy (anonymous):

one variable equations

OpenStudy (anonymous):

yeah then i think im doing things right unless you typed the equation wrong some how

OpenStudy (anonymous):

No i copy and pasted it

OpenStudy (anonymous):

aare you sure its -2 times -3 and not just -2 times 3

OpenStudy (anonymous):

yeah cuz if -2(x-3) the - follows the 3 when its -2(x+3) the 3 stays positive

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

also if you need more help there is this https://www.khanacademy.org/

OpenStudy (anonymous):

this is good too http://www.virtualnerd.com/

OpenStudy (anonymous):

Okay thanks! Do those websites do step by step? because if they don't then i'd be wasting my time

OpenStudy (anonymous):

yeah theres videos of them writing and talking about it

OpenStudy (anonymous):

Okay,Thanks for the help!!

OpenStudy (anonymous):

^ u ^ no problem

OpenStudy (anonymous):

o o U